# Dynamics of systems with varying number of particles: from Liouville equations to general master equations for open systems

### v1

@@ Abstract
A varying number of particles is the most relevant characteristic of systems of interest in nature and technology; ranging from the exchange of energy and matter with the surrounding environment to the change of particle numbers through internal dynamics such as reactions. The physico-mathematical modeling of these systems is extremely challenging, with the major difficulty being the time dependence of the number of degrees of freedom and the additional constraint that the increment or reduction of the number and species of particles must not violate basic physical laws. Theoretical models, in such a case, represent the key tool for the design of computational strategies for numerical studies that deliver trustful results. In this manuscript, we discuss complementary physico-mathematical approaches of varying numbers of particles inspired by different specific numerical goals and provide a generalized equation that embeds all of them.
@@ I Introduction
The conception of many-particle systems has played always a central role in modern theoretical and mathematical physics, from the many-electron systems of quantum mechanics to the tracers of fluid dynamics. Many-particle models are ubiquitous in condensed matter and provide the paradigm of reference for studying the properties of any existing substance. It is also an interesting approach per se in assessing the cascade of scales that characterize the physics of matter, since, in principle, it allows to systematically pass from the microscopic quantum mechanical resolution, up to the continuum hydrodynamics [3]. The models used in physics are in general based on an interaction potential, usually of a two-body nature, (e.g. electrostatic, Lennard-Jones) and the corresponding interlinked Newton’s equations of motion (or Schrödinger equations in quantum mechanics). This means that particle systems can be mathematically described as dynamical systems moving along calculable trajectories, which enables employing powerful mathematical and physical machinery to model, simulate and analyze them.
@@
Most of the machinery developed to model and simulate dynamical systems focuses on systems with a constant number of particles. However, this is not the case in many real-world applications, where we often have to focus on subsystems that are not isolated from material exchange with their surroundings. For instance, living cells constantly exchange molecules and energy with their environment; they consume chemical energy and dissipate heat. In terms of physical chemistry, every living system must be a nonequilibrium open system —a closed system has no life [40]. These nonequilibrium processes at the molecular scale often drive fundamental phenomena such as symmetry breaking, phase transitions and entropy production with profound impact on our understanding of living systems at meso- and macroscopic scales [43, 41, 37]. Moreover, the exchange of heat and matter is one of the main processes driving weather and climate phenomena, such as tipping events [49]. This manifests in the need for multiscale models [1], as well as reduced/coarse-grained models [35] and stochastic closure techniques [5]. Thus, to understand, model and simulate these processes, it is fundamental to develop mathematical and physical machinery to handle systems with varying particle numbers.
@@
From a mathematical perspective, it is a complex problem. The main difficulty is the on-the-fly change in the number of degrees of freedom of the system. At a differential equation level, the number of equations would change as the system changes particle numbers. The solution is to lift the dynamics to the space of densities/distributions, where the dynamics become linear, albeit infinite-dimensional and thus complicated to analyze. Some theoretical machinery has been developed in this direction [15, 24], but it often employs approaches based on quantum field theory, which are not a requirement to neither write nor understand the equations. From a computational perspective, one possible solution would be to consider a large enough closed subsystem that includes a substantial part of the environment. In this context, numerical methods of particle simulation have enormously risen in the past decades, improving our capability of performing precise calculations of large many-particle systems (see e.g. Refs.18, 6). However, despite the growing computational power, many systems of major interest are not yet affordable with standard available models, so we need alternative solutions that combine novel mathematical developments with computational approaches.
@@
In recent years, multiscale models and simulations have been developed to improve the efficiency of particle-based simulations. The basic idea is that one retains the degrees of freedom strictly required by the problem and simplifies at much coarser level the degrees of freedom that are not directly involved in the process of interest [39]. A prototype situation in most of the many-particle models is the necessity of coarse-graining the environment around the region of interest while retaining all the necessary details in the region itself; since the region of interest is subject to particle number fluctuations, the necessity of developing open many-particle systems rises naturally [44]. Moreover, once a mathematical model for the dynamics with varying numbers of particles is developed, it can naturally be adapted and applied to systems where the particle number fluctuations are provided not only by the external environment but also by the change of composition of a system, when for example in a mixture different species interact forming a third one. Despite the high dimensionality of the mathematical model, it can be used as a guiding framework to unify models at multiple scales and to derive physically consistent multiscale numerical schemes.
@@
In this work, we first introduce two frameworks to handle classical systems with varying particles (inspired by different physical problems and numerical approaches) and investigate the relations between them. The two main different points of view correspond to: (i) the approach based on Liouville-like equations of a subsystem embedded in a reservoir where the degrees of freedom are either explicitly integrated out, reducing the effect of the environment on the boundary conditions of the subsystem of interest[12, 31], or implicitly integrated out into a coupling term [34]; and (ii) a master equation where the diffusion process based on the Fokker-Planck operator is coupled with a reaction process that can change the number of particles. In this case, the particle description of the environment is not considered and its effect is modeled empirically a priori [11, 8]. The first approach is inspired by molecular simulations where the microscopic single molecular trajectories with explicit chemical details are accessible and thus the probability distribution function of the phase space of system and environment can be explicitly sampled [13]. This means that a simulation of an open system with the simplified environment obtained analytically by formal integration of its related degrees of freedom should deliver the same result as an equivalent subsystem in a simulation of the entire system; as a consequence, the validity of the theoretical model of the open system can be directly tested numerically (see e.g. Ref.[22]). The second approach is inspired by problems occurring at a larger scale than the microscopic molecular scale, e.g. processes at the scale of biological living cells, where the number of molecules is large enough to render molecular models intractable but not large enough to consider macroscopic approaches that neglect inherent stochastic fluctuations [42]. From a physics perspective, in this approach, molecules are represented as particles undergoing diffusion and chemical reactions are coarse-grained into events that simply change the chemical composition of the system. Thus, the effect of the environment is modeled empirically through the diffusion constant, the reaction rates and the stochastic effects. This model does not only provide a probabilistic model for reaction-diffusion processes [11, 8] but also serves as a starting point to derive other models at different scales. This yields a unifying theory that serves as the backdrop to derive numerical schemes that are consistent across multiple scales [33, 10].
@@
Independently from the original inspiration, both models deal with varying numbers of particles and thus must retain a common general structure within a unique framework. In this work, we proceed with the formal analysis of the similarities and differences between the two approaches. Based on this comparison, a generalized structure emerges in the form of a unitary equation, from which the previous approaches are special cases. Moreover, we discuss mathematical perspectives and potential problems of interest in physical and chemical applications, both from an analytical and numerical point of view.
@@
The paper is structured as follows. In section II, we explore two approaches to handle systems with Hamiltonian structure that are in contact with a material reservoir(s). Section III shows how to write the master equation to describe the probabilistic dynamics of reaction-diffusion processes and extends this to systems where the diffusion is governed by Langevin dynamics. In section IV, we further investigate the mathematical similarities between the previous approaches and use them to derive a general master equation for systems with varying particle numbers. To finalize, in section V, we discuss the perspectives and future scope of this work in the context of physical modeling and numerical simulations.
II Liouville-like equations for classical open systems
@@
The two approaches reported in this section were developed (the first one) or used (the second one) to conceptually frame a numerical method for the molecular simulation of open systems that exchange energy and matter with a large reservoir. The specific numerical code embedded in this theoretical framework is the Adaptive Resolution Simulation (AdResS) [13]; however, any molecular simulation approach that is characterized by the system-reservoir exchange of energy and particles could be framed as well in such models. The models presented here have been used as inspiration for designing and rationalizing the system-reservoir coupling [16, 32, 22]. Systems of molecular simulation are in general characterized by an explicit particle-particle Hamiltonian and by its corresponding phase-space probability distribution which leads to the calculation of statistical averages of physical quantities of interest [18]. The natural complete framework for the treatment of these systems is the Liouville equation, thus it seemed natural to manipulate the Liouville equation of the total system to obtain an equivalent equation for a subsystem where the surrounding (rest of the total system) has been explicitly integrated out in its particle degrees of freedom or, alternatively, the particle degrees of freedom have been empirically removed by modeling the reservoir as a generic thermodynamic bath.
II.1 From a large system of particles to an open subsystem of particles
@@
The model of open systems based on the Liouville equation of the total system of Ref 12 is here described in its essential features. Let us consider a large dynamical system of particles in equilibrium (called here Universe, ) and define an open subsystem, , containing particles (with the corresponding reservoir, , of particles, with ), as illustrated pictorially in Figure 1. Starting from the Liouville equations for the probability in phase space of the Universe and integrating out all the degrees of freedom of one would wish to derive a Liouville-like equation for the particles in taking into account that particles can freely move from the subsystem to the reservoir and vice versa.
@@
The integration of Eq. 4 w.r.t. the variables of implies a straightforward calculation procedure which does not carry any relevant conceptual aspects and thus it is not reported here; the corresponding details of can be found in Refs. 12, 32, here the final results are reported:
@@
the latter corresponds to the mean field force that the outer particles exert onto the th particle in . is defined as the conditional distribution of an outer particle at the given state of an inner particle; one can assume that it is known or a model can be proposed. Furthermore,
@@
In Eq. 13 corresponds to the (modeler’s) assumption of the reservoir’s one particle distribution calculated at the interface boundary . It must be also added that the addition of an external thermostat in the reservoir does not change the equation and allows to extend the model to situations of non-equilibrium like thermal gradient [31].In general, such a theoretical framework rationalizes the numerical scheme AdResS; the idea is explicitly discussed in the section below.
II.1.1 Theoretical scheme and AdResS method
The Adaptive Resolution (AdResS) scheme for molecular dynamics is a numerical approach where a region of interest is treated at high molecular resolution while the surrounding is treated at a coarser level. Molecules can freely move from one region to the other and accordingly change their resolution [38]. The model of open systems discussed above (section II.1) allows a further simplification of the environment around the system of interest. In fact, the molecular resolution outside the system of interest can be drastically reduced to non-interacting point-particles embedded in a mean-field [13]; the properties of this mean-field are defined by the coupling terms, and , of the model discussed in section II.1.
Below we report the basic idea, while Fig.2 illustrates the numerical scheme and its correspondence to the theoretical model.
The system-reservoir coupling terms of the model discussed in II.1, are:
and
All the physical quantities involved in eqs. 14 and 15 are directly accessible in a simulation (i.e. can be explicitly calculated in ) except for and which, lacking molecular resolution in the reservoir, cannot be calculated and need to be modeled/assumed. For physical consistency, such quantities must correspond to those calculated in a large molecular system that contains the open system as a subsystem. In terms of the algorithm, the conditions from eqs. 14 and 15 imply the assumption that, at the interface, and are the ones expected in a simulation at full molecular resolution. To this aim, in AdResS, at the interface region has been imposed a so-called thermodynamic force, , derived self-consistently in the simulation scheme during the equilibration stage (see also [19]). This enforces a particle density (i.e. ) equivalent to the particle density of the reference fully resolved system together with a thermostat that fixes the temperature at the target value. Thus, is the mean-field one-particle potential in which the molecules at the interface and in the reservoir are embedded. For , it has been found that fixing , for molecular systems of practical interest, leads to a two-body distribution function (i.e. ), which closely reproduces the two-body distribution function of the reference fully resolved system. However, in case of necessity, a two-body thermodynamic force, in the same fashion of the one-body force that fixes , can be imposed as well [47]. Once one can be assured that and in are as expected, it follows that the coupling terms of eq. 14 and eq. 15 are implicitly realized in the simulation since all the other quantities can be explicitly calculated. This means the set up of the simulation contains all the information to reproduce the physics expressed by the Liouville-like equation for . It must be mentioned that the model of section II.1 is valid also for situations out of equilibrium, e.g. a thermal or density gradient imposed on the system by a fluctuating reservoir or by multiple reservoirs at different thermodynamic conditions. Such features have been incorporated in the AdResS scheme and were numerically tested with satisfactory results [32, 22].
II.2 The Bergman-Lebowitz equation of open systems

### v2

@@ Abstract
A varying number of particles is one of the most relevant characteristics of systems of interest in nature and technology, ranging from the exchange of energy and matter with the surrounding environment to the change of particle number through internal dynamics such as reactions. The physico-mathematical modeling of these systems is extremely challenging, with the major difficulty being the time dependence of the number of degrees of freedom and the additional constraint that the increment or reduction of the number and species of particles must not violate basic physical laws. Theoretical models, in such a case, represent the key tool for the design of computational strategies for numerical studies that deliver trustful results. In this manuscript, we review complementary physico-mathematical approaches of varying number of particles inspired by rather different specific numerical goals. As a result of the analysis on the underlying common structure of these models, we propose a unifying master equation for general dynamical systems with varying number of particles. This equation embeds all the previous models and can potentially model a much larger range of complex systems, ranging from molecular to social agent-based dynamics.
@@ I Introduction
The conception of many-particle (or equivalently, many-body) systems has played a central role in modern theoretical and mathematical physics, from the many-electron systems of quantum mechanics to the tracers of fluid dynamics. Many-particle models are ubiquitous in condensed matter and provide the paradigm of reference for studying the properties of any existing substance. It is also an interesting approach per se in assessing the cascade of scales that characterize the physics of matter, since, in principle, it allows to systematically pass from the microscopic quantum mechanical resolution, up to the continuum hydrodynamics [3]. The models used in physics are in general based on an interaction potential, usually of a two-body nature, (e.g. electrostatic, Lennard-Jones) and the corresponding interlinked Newton’s equations of motion (or Schrödinger equations in quantum mechanics). This means that particle systems can be mathematically described as dynamical systems moving along calculable trajectories, enabling powerful mathematical and physical machinery to model, simulate and analyze them.
@@
Most of the machinery developed to model and simulate molecular or particle-based dynamical systems focuses on systems with a constant number of particles, this means they are closed but not necessarily isolated from heat exchange. However, this is not the case in many real-world applications, where we often have to focus on subsystems that exchange material with their surroundings; we refer to such systems as open systems. For instance, living cells constantly exchange molecules and energy with their environment; they consume chemical energy and dissipate heat. In terms of physical chemistry, every living system must be a nonequilibrium open system —a closed system has no life [62]. These nonequilibrium processes at the molecular scale often drive fundamental phenomena such as symmetry breaking, phase transitions and entropy production with profound impact on our understanding of living systems at meso- and macroscopic scales [65, 63, 57]. Moreover, the exchange of heat and matter is one of the main processes driving weather and climate phenomena, such as tipping events [75]. This manifests in the need for multiscale models [1], as well as reduced/coarse-grained models [53] and stochastic closure techniques [6]. Thus, to understand, model and simulate these processes, it is fundamental to develop mathematical and physical machinery to handle systems with varying particle number.
@@
From a mathematical perspective, it is a complex problem. The main difficulty is the on-the-fly change in the number of degrees of freedom of the system. At a differential equation level, the number of equations would change as the system changes particle number. The solution is to lift the dynamics to the space of densities/distributions, where the dynamics become linear, albeit infinite-dimensional and thus complicated to analyze. This is analogous to switching representation from a nonlinear dynamical system to the linear Liouville equation, or in the case of stochastic dynamical systems to a Fokker-Planck equation. However, note some information is lost in the process, as we can no longer track individual particles. Following this reasoning, some theoretical machinery has been developed in previous works to handle systems with varying particle number. We will overview the main ones in this manuscript. Other similar approaches often employ methods based on quantum field theory [24, 35, 14, 15], which are not a requirement to neither write nor understand the equations. From a computational perspective, one possible solution would be to consider a large enough closed subsystem that includes a substantial part of the environment. In this context, numerical methods of particle simulation have enormously risen in the past decades, improving our capability of performing precise calculations of large many-particle systems (see e.g. [28, 7]). However, despite the growing computational power, many systems of major interest are not yet affordable with standard available models, so we need alternative solutions that combine novel mathematical developments with computational approaches.
@@
In recent years, multiscale models and simulations have been developed to improve the efficiency of particle-based simulations. The basic idea is that one retains the degrees of freedom strictly required by the problem and simplifies the degrees of freedom that are not directly involved in the process of interest at coarser levels [60]. A prototype situation in most of the many-particle models is the necessity of coarse-graining the environment around the region of interest while retaining all the necessary details in the region itself; since the region of interest is subject to particle number fluctuations, the necessity of developing open many-particle systems rises naturally [66]. Moreover, once a mathematical model for the dynamics with varying number of particles is developed, it can organically be adapted and applied to systems where the particle number fluctuations are provided not only by the external environment but also by the change of composition of a system, when for example in a mixture different species interact forming a third one. Despite the high dimensionality of the resulting mathematical model, it can be used as a guiding framework to unify models at multiple scales and to derive physically consistent multiscale numerical schemes.
The concept of subsystem is closely related to the concept of subdynamics for which there is a vast literature available. For example, the model proposed by Prigogine and coworkers [61, 31], the model of Emch and Sewell [26], and the model of Robertson[67, 69], to cite but a few. The idea is based on the projection method of Zwanzig [77] for quantum systems (used also for classical systems). The evolution of the probability density of the system in the von Neumann or Liouville equation is projected on a subspace with a reduced number of variables and the rest of the system is formally coarse-grained. The models discussed in this paper share the idea of integrating/coarse-graining the degrees of freedom outside the subsystem, however they add the explicit treatment of the number of particles as a variable of the problem. This characteristic is not explicitly mentioned in the literature cited above, thus leaving the impression that the number of particles of the subsystem is assumed to be fixed. In a system with varying number of particles one should explicitly discuss the normalization of the probability density of the subsystem; this operation is substantially different from the case of a fixed number of particles. In addition, the equations proposed in [61, 31, 26, 67, 69], while mathematically rigorous and certainly elegant, are characterized by a rather complex structure difficult to implement in modern computational simulation techniques. Instead, for computational implementation, one needs models that are characterized by a certain level of mathematical rigor but at the same time are constructed to efficiently capture the relevant physics. The model of Bergman and Lebowitz [51, 52], discussed later on, can be considered the first historical attempt towards such an effort and became a source of inspiration for most of the progress reported here.
@@
In this work, we first review two frameworks to handle classical systems with varying number of particles (motivated by different physical problems and numerical approaches). We introduce these frameworks and investigate the relations between them to show how a general master equation for systems varying number of particles emerges. The two main different points of view correspond to: (i) the approach based on Liouville-like equations of a subsystem embedded in a reservoir where the degrees of freedom are either explicitly integrated out, reducing the effect of the environment on the boundary conditions of the subsystem of interest [20, 44], or implicitly integrated out into a coupling term [51]; and (ii) a master equation where the diffusion process based on the Fokker-Planck operator is coupled with a reaction process that can change the number of particles. In this case, the particle description of the environment is not considered and its effect is modeled empirically a priori [18, 14]. The first approach is inspired by molecular simulations where the microscopic single molecular trajectories with explicit chemical details are accessible and thus the probability distribution function of the phase space of system and environment can be explicitly sampled [21]. This means that a simulation of an open system with the simplified environment obtained analytically by formal integration of its related degrees of freedom should deliver the same result as an equivalent subsystem in a simulation of the entire system; as a consequence, the validity of the theoretical model of the open system can be directly tested numerically (see e.g. [33]). The second approach is inspired by problems occurring at a larger scale than the microscopic molecular scale, e.g. processes at the scale of biological living cells, where the number of molecules is large enough to render molecular models intractable but not large enough to consider macroscopic approaches that neglect inherent stochastic fluctuations [64]. From a physics perspective, in this approach, molecules are represented as particles undergoing diffusion and chemical reactions are coarse-grained into events that simply change the chemical composition of the system. Thus, the effect of the environment is modeled empirically through the diffusion constant, the reaction rates and the stochastic effects. This model does not only provide a probabilistic model for reaction-diffusion processes [18, 14] but also serves as a starting point to derive other models at different scales. This yields a unifying theory that serves as the backdrop to derive numerical schemes that are consistent across multiple scales [49, 17, 12].
@@
Independently from the original inspiration, both models deal with varying number of particles and thus must retain a common general structure within a unique framework. In this work, we proceed with the formal analysis of the similarities and differences between the two approaches. Based on this comparison, a generalized structure emerges in the form of a unitary equation, from which the previous approaches are special cases. This generalized structure can be used for any dynamical system in an open setting, enabling a broader range of applications within physics and beyond. To finalize, we discuss mathematical perspectives and potential problems of interest in physical and chemical applications both from an analytical and numerical point of view. We further discuss possible applications of the general master equation beyond the fields that initially motivated the equation.
@@
The paper is structured as follows. To motivate the reader, section II focuses on an intuitive overview of the final general master equation for systems with varying particle number, which emerges naturally in section V. However, the equation itself is originally motivated by the specific approaches presented in sections III and IV. More specifically, in section III, we explore two approaches to handle systems with Hamiltonian structure that are in contact with a material reservoir(s). In section IV, we show how to write the master equation to describe the probabilistic dynamics of reaction-diffusion processes, and we extend this to systems where the diffusion is governed by Langevin dynamics. In section V, we show how the mathematical similarities between the approaches from sections III and IV inspire formulating the general master equation and how it recovers the previous approaches as special cases. Finally, in section VI, we discuss the perspectives and future scope of this work in the context of physical modeling, numerical simulations and applications in other fields.
II A general master equation for systems with varying number of particles
The general master equation presented in this section was inspired by the frameworks presented in sections III and IV. For educational purposes, we first carry out an intuitive derivation of the equation. We formulate the master equation for systems with varying number of particles starting from a general random dynamical system perspective. This together with section V constitute the main results of the paper.
We first consider a random dynamical system with fixed dimension , e.g. a system with a fixed number of particles. The dynamics of the trajectories of the systems are given by the system of stochastic differential equations
where could represent many things, depending on the application at hand. In physics, it could represent the position of particles or perhaps the positions and velocities of particles. In other fields like social systems, it could represent the position of an agent in opinion space [37] or the population of individuals of a certain species [50]. The function incorporates the deterministic drift as well as noise components. One can equivalently write the dynamics of the probability distribution of the system in phase space given by the corresponding Fokker-Planck equation
with the infinitesimal generator of the Fokker-Planck equation for the corresponding particle system, and the probability density of the system being at at time , which naturally integrates to one when integrating over the whole phase space. Reasonable boundary conditions for are usually reflective in a bounded domain or vanishing at infinity for a non-bounded domain.
To write a similar description for a system with a variable number of particles, one cannot write it in the form of eq. 1 since the dimension of the system changes as time evolves. However, we can write something similar to the master equation from eq. 2. Consider the family of distributions , where is the probability density of having particles (or whatever else we are modeling) at positions or states at time . The phase space now has a much more complex structure with continuous and discrete degrees of freedom, so the normalization condition is
where the integral is over the whole available space. The master equation for the system with varying particle number can then be written as a family of Fokker-Planck equations, one for each -particle level. These are then coupled by operators that transfer probability from the -particle state to the one modeling processes that change the particle number (e.g. interactions with a material reservoir or reactions). These operators form a coupling matrix, which incorporates into the master equation
Each row of the equation corresponds to the dynamics of the -particle distribution. If the coupling matrix is removed, one obtains a set of uncoupled Fokker-Planck equations, each for a different particle number. If the space dependence is removed, one recovers a continuous-time Markov chain, where the operators become transition rates. This master equation for systems with varying particle number can be written more compactly by focusing on the th component
with . Although these infinite sums might seem problematic at first, in practical cases most are zero and the sums remain finite, as we will see later. Moreover, in analogy with stochastic matrices in continuous-time Markov chains, we expect each matrix column to sum to zero to enforce the conservation of probability. As in this case, we have a matrix of operators and space dependence, the probability conserving condition translates into the following constraint imposed into the coupling operators
for any given test arbitrary probability density for particles and where the integrals run over the whole available space. Each of these integrals represents the probability flux associated with a transition. By forcing their sum to be zero, we enforce that the net probability flux leaving one state is the same as the one entering another state(s). Thus, enforcing probability conservation. As a final note, there is a connection between this description and hybrid switching diffusions [56, 76, 8, 13]. However, in the particle context, hybrid switching diffusions can only account for a change in the state of a particle but cannot incorporate or remove particles.
In the next sections we will review the derivation of master equations of this form written for specific contexts, which have inspired the generalization above that applies to a larger class of systems with varying number of particles [24, 20, 44, 14, 18, 52]. In sections III and IV, we will explore specific physical systems that involve varying particle number. Then, in section V, we will show how these specific systems inspired the general master equation (eq. 5) and how they are recovered as particular cases.
III Liouville-like equations for classical open systems
@@
The two approaches reported in this section were developed (the first one) or used (the second one) to conceptually frame a numerical method for the molecular simulation of open systems that exchange energy and matter with a large reservoir. The specific numerical code embedded in this theoretical framework is the Adaptive Resolution Simulation (AdResS) [29, 71, 21]; however, any molecular simulation approach that is characterized by the system-reservoir exchange of energy and particles could be framed as well in such models. The models presented here have been used as inspiration for designing and rationalizing the system-reservoir coupling [25, 45, 33]. Systems of molecular simulation are in general characterized by an explicit particle-particle Hamiltonian and by its corresponding phase-space probability density. This latter is not known explicitly, however it is statistically sampled either through a single long trajectory or through a collection of short trajectories, each with an initial condition uncorrelated to the initial condition of the others. As a consequence the statistical calculation of physical quantities is done by sampling and averaging the physical quantity of interest along such trajectories [28]. The natural complete framework for the treatment of these systems is the Liouville equation, thus it seemed natural to manipulate the Liouville equation of the total system to obtain an equivalent equation for a subsystem where the surrounding (rest of the total system) has been explicitly integrated out in its particle degrees of freedom or, alternatively, the particle degrees of freedom have been empirically removed by modeling the reservoir as a generic thermodynamic bath.
III.1 From a large system of particles to an open subsystem of particles
@@
The model of open system based on the Liouville equation of the total system of [20] is here described in its essential features. Let us consider a large dynamical system of particles in equilibrium (called here Universe, ) and define an open subsystem, , containing particles (with the corresponding reservoir, , of particles, with ), as illustrated pictorially in fig. 1. Starting from the Liouville equations for the probability in phase space of the Universe and integrating out all the degrees of freedom of one would wish to derive a Liouville-like equation for the particles in taking into account that particles can freely move from the subsystem to the reservoir and vice versa.
@@
The integration of eq. 10 w.r.t. the variables of implies a straightforward calculation procedure which does not carry any relevant conceptual aspects and thus it is not reported here; the corresponding details of can be found in [20, 45], here the final results are reported:
@@
The latter corresponds to the mean field force that the outer particles exert onto the th particle in . The quantity is defined as the conditional distribution of an outer particle at the given state of an inner particle; one can assume that it is known or a model can be proposed. Furthermore,
@@
In eq. 19 corresponds to the (modeler’s) assumption of the reservoir’s one particle distribution calculated at the interface boundary . It must be also added that the addition of an external thermostat in the reservoir does not change the equation and allows us to extend the model to situations of non-equilibrium like thermal gradient [44]. In general, such a theoretical framework rationalizes the AdResS molecular dynamics protocol, but in principle, as said above, it can be applied to any other molecular simulation method that aims at considering a varying number of molecules.
III.2 The Bergman-Lebowitz equation of open systems